OpenStudy (anonymous):

The rate of water flow R(t) can be estimated by W(t) = ln(t2 + 7). Use W(t) to approximate the average rate of water flow during the 8-hour time period. Indicate units of measure.

OpenStudy (anonymous):

@freckles @IrishBoy123 @zepdrix I would integrate correct? If so how would I integrate ln?

OpenStudy (freckles):

do you know integration by parts?

OpenStudy (freckles):

also what 8 hour time period are they talking about?

OpenStudy (anonymous):

Leme give you a screenshot of that sorry. And I am familiar with the term but it would be nice if you could show me how it's done.

OpenStudy (anonymous):

https://gyazo.com/96debaa25dbffb8f00cf7b37bdf6365b

OpenStudy (anonymous):

Would u-substitution not work?

OpenStudy (freckles):

you will have to use integration by parts if you want to integrate W though your less has been all about approximating integrals using midpoint, trapezoid, and other rules so you may can get away without doing integration by parts.... however if integration by parts is the way you want to choose to go... then take this as an example \[\int\limits 1 \cdot \ln(x) dx\] notice the times 1 there we need the 1 1 will be the one we choose to integrate and ln(x) will be the one we differentiate \[\int\limits \color{red}{1} \cdot \ln(x) dx\\ =\color{red}{x} \color{blue}{\ln(x)}- \int\limits \color{red}{x} \cdot \color{blue}{\frac{1}{x}} dx\]

OpenStudy (freckles):

\[\int\limits \ln(x) dx=x \ln(x)-\int\limits 1 dx \\ =x \ln(x)-x +C\]

OpenStudy (freckles):

you will do something similar here

OpenStudy (freckles):

you have you want to evaluate \[\frac{ \int\limits _0^8 W(t) dt}{8-0}\]

OpenStudy (freckles):

\[\int\limits W(t) dt= \int\limits 1 \cdot \ln(t^2+7) dt\] integrate 1 and differentiate t^2+7 you can plug into this formula: \[\int\limits f'(x) g(x) dx=f(x) g(x)-\int\limits f(x) g'(x)dx\] notice I chose f' to integrate and g to differentiate

OpenStudy (freckles):

you will have a bit more work to do like a trig identity unless you know your integration table well

OpenStudy (freckles):

at the end

OpenStudy (freckles):

now you might decide to approximate that integral above using some method you have learned in this I guess its called lesson pack

OpenStudy (freckles):

integrate 1 and differentiate ln(t^2+7) * working in the dark

OpenStudy (anonymous):

Can you show how it applies to ln? It'll help grasp it.

OpenStudy (freckles):

how what applies to ln?

OpenStudy (anonymous):

Integration by parts.

OpenStudy (freckles):

oh you didn't see the example?

OpenStudy (freckles):

\[\int\limits f'(x)g(x)dx=f(x) g(x)-\int\limits f(x)g'(x) dx \\ \int\limits 1 \cdot \ln(x) dx= x \ln(x)-\int\limits x \frac{1}{x} dx \\ \int\limits 1 \cdot \ln(x) dx=x \ln(x)-\int\limits 1 dx=x \ln(x)-x+C \\ \text{ so } \int\limits \ln(x) dx=x \ln(x)-x+C\] this example above

OpenStudy (freckles):

you have \[\int\limits 1 \cdot \ln(x^2+7) dx\]

OpenStudy (freckles):

choose f'(x)=1 and the other function, the log function, to be g(x) just like I did in the example

OpenStudy (freckles):

have you plug into the integration by parts formula above?

OpenStudy (freckles):

f(x)=x and g'(x)=? if g(x)=ln(x^2+7)

OpenStudy (anonymous):

Absolutely confused on how I should approach the part with ln(t^2+7)

OpenStudy (anonymous):

Just wanted to mention you can use a substitution, but it doesn't prevent you from having to use integration by parts. Set \(t=\sqrt{e^s-7}\), so \(\mathrm{d}t=\dfrac{e^s}{2\sqrt{e^s-7}}\,\mathrm{d}s\), and \(s=\ln(t^2+7)\). This means you get \[\int_0^8\ln(t^2+7)\,\mathrm{d}t=\frac{1}{2}\int_{\ln7}^{\ln71}\frac{se^s}{\sqrt{e^s-7}}\,\mathrm{d}s\]This way, you end up making more work for yourself. Instead, using freckles' suggestion, you have \[\int\ln(t^2+7)\,\mathrm{d}t=uv-\int v\,\mathrm{d}u\]where \[\begin{matrix}u=\ln(t^2+7)&&&\mathrm{d}v=\mathrm{d}t\\[1ex] \mathrm{d}u=\dfrac{2t}{t^2+7}\,\mathrm{d}t&&&v=t\end{matrix}\]so \[\int_0^8\ln(t^2+7)\,\mathrm{d}t=\bigg[t\ln(t^2+7)\bigg]_0^8-2\int_0^8\frac{t^2}{t^2+7}\,\mathrm{d}t\]

OpenStudy (anonymous):

I'm probably going to need to practice this a lot more barely got it. Thanks for your help guys.